OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (anonymous):

umm, take the first equation and isolate y

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

\[3x=6y+3\\ x=2y+1\\ 7(2y+1)-5y=-11\\ 14y+7-5y=-11\\ 9y=-18\\ y=-2\\ 3x+12=3\\ x=-3\]